ELECTRON ATTACHMENT AND DETACHMENT IN THE WATER HEXAMER ANION

\underline{CHRISTOPHER F. WILLIAMS}, and JOHN M. HERBERT, Department of Chemistry, Ohio State University, 100 W. 18th Ave. Columbus, OH 43210.

AIMD simulations have been performed on isomers of the water hexamer anion at temperatures comparable to those found in molecular beams. Electron detachment from these species has been simulated using a sudden approximation where the anion geometry at a selected time is propagated using ordinary MD with a neutral water potential. The resulting distribution of conformers and the relationship between the electron attachment and detachment pathways will be discussed.